A ladder is leaning against a building so that the distance from the ground to the top of the ladder is 1 foot less than the length of the ladder. Find the length of the ladder if the distance from the bottom of the ladder to the building is 3 feet.
Answer by stanbon(75887)

A ladder is leaning against a building so that the distance from the ground to the top of the ladder is 1 foot less than the length of the ladder. Find the length of the ladder if the distance from the bottom of the ladder to the building is 3 feet.
----
Sketch the picture.
You have a right triangle with hypotenuse = h
base = 3
height = h-1
-------
Equation:
h^2 = 3^2 + (h-1)^2
h^2 = 9 + h^2 -2h + 1
-----
2h = 10
height = h-1 = 4 ft.
ladder = h = 5 ft.
-------------
Cheers,
Stan H.
----------------